Attendance at Friday's soccer game was 792. This was 2 3/4 times as many fans as at Thursday's game. What was the attendance for Thursday's game?
Answer by harpazo(655)


Let's begin by writing your mixed fraction as an improper fraction. So,
2_3/4 can be written as 11/4.
Let x = the attendance on Thursday.
We know that 792 people attended Friday's game.
Here is your equation:
(11/4)(x) = 792, which can also be written as (11x/4) = 792
Multiply both sides by 4 to remove the fraction on the left side.
4 * (11x/4) = 792 * 4
11x = 3168
Divide 3168 by 11 to find x.
x = 3168/11
x = 288
So, 288 people attended Thursday's game.